1/3 of a certain number added to 1/5 of the next consecutive number is 5.what is the sum of the 2 numbers?
Answer by lwsshak3(11628)

1/3 of a certain number added to 1/5 of the next consecutive number is 5.what is the sum of the 2 numbers?
**
let x=a certain number
x+1=next consecutive number
x/3+(x+1)/5=5
LCD:15
5x+3x+3=75
8x=72
x=9
x+1=10
sum of the 2 numbers=9+10=19